# Thermodynamic and quantum fluctuations of horizon area

### v3

@@ II Quantum fluctuations of black hole entropy
The coefficient $4$ in the area variance in Eq.(4) differs from the corresponding coefficients in Refs. Ciambelli2025 ; ParikhPereira2024 . However, the same equation (4) can be obtained from the thermodynamic fluctuations of the horizon entropy in the de Sitter (dS) spacetime, see next Section.
@@ III Thermodynamic fluctuations of de Sitter horizon entropy
Here $H$ is the Hubble parameter and $T=H/\pi$ is the local temperature of the de Sitter state. This local temperature is twice the Gibbons-Hawking temperature $T_{\rm GH}=H/2\pi$ associated with the cosmological horizon. Nevertheless, the total entropy $S=sV_{H}$ of the Hubble volume $V_{H}=(4\pi/3)/H^{3}$ coincides with the Gibbons-Hawking entropy of cosmological horizon, $sV_{H}=A/4G$. This demonstrates, that the local entropy obeys the holographic bulk-horizon correspondence.Hooft1993 ; Susskind1995 This holographic correspondence connects the extensive entropy of the homogeneous Universe with the non-extensive entropy of the cosmological horizon.
The factor 2 has a natural explanation from the point of view of two observers. The observer measuring the local temperature, such as the activation temperature of the ionization of atoms in de Sitter environment, has all the necessary information. On the other hand, the observer measuring the temperature of the Hawking radiation from the cosmological horizon has no information about the simultaneous creation of the Hawking partner beyond the horizon. This is the reason why this observer underestimates the temperature by a factor of 2. These arguments do not apply to a black hole, where the creation of a Hawking partner is represented by a back reaction.Wilczek2000
Since the de Sitter state is homogeneous and isotropic, we can apply the standard equations for the thermodynamic fluctuations (see the book Landau_Lifshitz by Landau and Lifshitz). Then using the linear dependence of local entropy on temperature and thus the linear dependence of the heat capacity, one obtains the fluctuations of the total entropy in the Hubble volume $V_{H}$:
@@
The Eq.(6) does not contain $\hbar$, which demonstrates that it describes thermodynamic fluctuations. However, since $S=A/4G\hbar$, this is equivalent to the Eq.(4) describing quantum fluctuations. The exact relation between the quantum and thermodynamic fluctuations is typical for the systems with horizons. This can be seen also from the processes of splitting of the black hole into the smaller parts. On one hand, such process is quantum, being determined by the macroscopic quantum tunneling, and on the other hand it is determined by the decrease of entropy after splitting,Wilczek2000 ; Volovik2022 which is manifestation of the thermal fluctuations. This is one of many examples where gravity serves as a bridge between thermodynamics and quantum mechanics.Wald1999
@@ IV Black hole fluctuations via ensemble of Planck scale ”atoms”
Another interesting property of the quantum-thermodynamic fluctuations of the black hole is provided by the toy model of the black hole entropy.Volovik2025P In this model, the black hole with mass $M$ is represented by the peculiar ensemble of $N$ Planckons – objects with reduced Planck massMarkov1967 ; Hawking1971 ; Aharonov1987 ; Rovelli2024 ; YenChinOng2025 ; Trivedi2025 :
$M=Nm_{\rm P}\,\,,\,\,m_{\rm P}^{2}=\frac{K}{2\pi}=\frac{1}{8\pi G}\,.$ (8)
@@
In the thermodynamic limit $N\gg 1$, the equation (6) gives for fluctuations of the Planckon number:
@@
This non-extensive composition law is in the origin of the black-hole thermodynamics, which distinguishes the Tsallis-Cirto $\delta=2$ entropy from other possible types of generalized entropy.Odintsov2022a ; Odintsov2022b ; Odintsov2025
On the other hand, the ensemble of the $\cal N$} pairs of Planckons has conventional composition law $S({\cal N}_{1}+{\cal N}_{2})=S({\cal N}_{1})+S({\cal N}_{2})$ and conventional fluctuations, which are consistent with the Poisson distribution:
@@ V Discussion
(ii) The variance of the entropy of the cosmological horizon in de Sitter state can be obtained in the domain of classical thermodynamics fluctuations. The entropy of the horizon is equal to the entropy of the Hubble volume, thus representing the holographic bulk-horizon connection. As a result, the variance of the entropy can be obtained from the conventional thermodynamics in bulk, which determines thermal fluctiations of thermodynamic variables, such as entropy. The result obtained in the domain of thermal fluctuations coincides with the result obtained for black hole in the domain of quantum fluctuations in Eq.(4).

### v4

@@ II Quantum fluctuations of black hole entropy
The coefficient $4$ in the area variance in Eq.(4) differs from the corresponding coefficients in Refs. Ciambelli2025 ; ParikhPereira2024 . In Ref.Ciambelli2025 the coefficient is $\pi$, while in Ref. ParikhPereira2024 it was concluded that the precise relationship between the computed quantum variance and fluctuations associated with black hole entropy remains unclear. However, the same equation (4) can be obtained for the variance of the area of the cosmological horizon. The latter can be obtained by considering thermodynamic fluctuations of the entropy of the de Sitter (dS) spacetime, see the next Section.
@@ III Thermodynamic fluctuations of de Sitter horizon entropy
Here $H$ is the Hubble parameter and $T=H/\pi$ is the local temperature of the de Sitter state, which in particular determines the ionization rate of an atom in de Sitter environment. This local temperature is twice the Gibbons-Hawking temperature $T_{\rm GH}=H/2\pi$ associated with the cosmological horizon. Nevertheless, the total entropy $S(V_{H})=sV_{H}$ of the Hubble volume $V_{H}=(4\pi/3)/H^{3}$ coincides with the Gibbons-Hawking entropy of cosmological horizon, $S(V_{H})=A/4G$. This demonstrates, that the local entropy obeys the holographic bulk-horizon correspondence.Hooft1993 ; Susskind1995 This holographic correspondence connects the extensive entropy of the homogeneous Universe with the non-extensive entropy of the cosmological horizon.
The difference between the local temperature and the Gibbons-Hawking temperature has a natural explanation: these temperatures are measured by two different observers. The observer measuring the local temperature, such as the activation temperature of the ionization of atoms in de Sitter environment, has all the necessary information. On the other hand, the observer measuring the temperature of the Hawking radiation from the cosmological horizon has no information about the simultaneous creation of the Hawking partner beyond the horizon. This is the reason why this observer underestimates the temperature by a factor of 2. These arguments do not apply to a black hole, where the creation of a Hawking partner is represented by a back reaction.Wilczek2000
Since the de Sitter state is homogeneous and isotropic, we can apply the standard equations for the thermodynamic fluctuations (see Sec. 112 in the book Landau_Lifshitz by Landau and Lifshitz). Then using the linear dependence of local entropy density on temperature and thus the linear dependence of the heat capacity, one obtains the fluctuations of the local entropy:
$\left<(\Delta s)^{2}\right>=\frac{s}{V}\,.$ (6)
Then for the total entropy $S(V)=sV$ in the arbitrary volume $V$ one has:
@@
The Eq.(7) does not contain the Planck constant $\hbar$, which demonstrates that it describes thermodynamic fluctuations. This equation is automatically valid also for the entropy of the Hubble volume with $V=V_{H}$. Then, since $S(V_{H})=sV_{H}=A/4G\hbar$, one obtains for the fluctuations of the area of the cosmological horizon:
$\left<(\Delta A)^{2}\right>=4G\hbar\left<A\right$>\,. (8)
This coincides with Eq.(4) for the quantum fluctuations of the area of the black hole horizon. Thus the classical fluctuations of the entropy of the Hubble volume are equivalent to the quantum fluctuations of the area of the cosmological horizon.
The exact relation between the quantum and thermodynamic fluctuations is typical for the systems with horizons. This can be seen also from the processes of splitting of the black hole into the smaller parts. On one hand, such process is quantum, being determined by the macroscopic quantum tunneling, and on the other hand it is determined by the decrease of entropy after splitting,Wilczek2000 ; Volovik2022 which is manifestation of the thermal fluctuations. This is one of many examples where gravity serves as a bridge between thermodynamics and quantum mechanics.Wald1999
@@
Finally, note that we consider the ideal de Sitter state, ignoring its instability and decay.Volovik2024T The latter disrupt the connection between the properties of the cosmological and black hole horizons, see also Refs.Trivedi2025a ; Trivedi2025b .
@@ IV Black hole fluctuations via ensemble of Planck scale ”atoms”
Another interesting property of the quantum-thermodynamic fluctuations of the black hole is provided by the toy model of the black hole entropy.Volovik2025P In this model, the black hole with mass $M$ is represented by the peculiar ensemble of $N$ Planckons – objects with Planck scale mass.Markov1967 ; Hawking1971 ; Aharonov1987 ; Rovelli2024 ; YenChinOng2025 ; Trivedi2025 . For Planckons with reduced Planck mass $m_{\rm P}$:
$M=Nm_{\rm P}\,\,,\,\,m_{\rm P}^{2}=\frac{\hbar}{8\pi G}\,.$ (10)
@@
In the thermodynamic limit $N\gg 1$, which according to Eq.(10) also corresponds to the classical limit $\hbar\rightarrow 0$, the equation (7) gives for fluctuations of the Planckon number:
@@
This non-extensive composition law is in the origin of the black-hole thermodynamics, which distinguishes the Tsallis-Cirto $\delta=2$ entropy from other possible types of generalized entropy.Odintsov2022a ; Odintsov2022b ; Odintsov2025 The same composition law for the entropies of the inner and outer horizons of the rotating Kerr black hole gives the following entropy of the Kerr black hole:Volovik2025Kerr
$S_{\rm Kerr}(N,J)=4\pi N(N-1)+4\pi\sqrt{J(J+1)}\,$, (14)
where the Plackons with Planck mass $M_{\rm P}=\sqrt{\hbar/G}$ were used.
On the other hand, the ensemble of the $\cal N$} pairs of Planckons has conventional extensive composition law:
$S({\cal N}_{1}+{\cal N}_{2})=S({\cal N}_{1})+S({\cal N}_{2})\,$, (15)
and the conventional fluctuations, which are consistent with the Poisson distribution:
@@ V Discussion
(ii) The variance of the entropy of the cosmological horizon in de Sitter state can be obtained in the domain of classical thermodynamics fluctuations. The entropy of the cosmological horizon is equal to the entropy of the Hubble volume, thus representing the holographic bulk-horizon connection. As a result, the variance of the entropy can be obtained from the conventional thermodynamics in bulk, which determines thermal fluctuations of thermodynamic variables, such as entropy. The variance of the entropy in Eq.(7), which is obtained in the domain of thermal fluctuations, is applicable to the arbitrary volume $V$ of the homogeneous de Sitter state. When applied to the Hubble volume $V=V_{H}$, one obtains the variance of the cosmological horizon area in Eq.(8) in agreement with the equation (4), which was obtained for the black hole horizon in the domain of quantum fluctuations.